OpenStudy (anonymous):

v+5 v --- + --- add and simplify if possible v v+5

OpenStudy (dumbcow):

common denominator is v(v+5)

OpenStudy (sriram):

as said above common denominator=v(v+5) first fraction can be written as (v+5) (v+5) (v+5)^2 ---- x ---- = ------- v (v+5) v(v+5) and similarly the second fraction as v^2 ---- (v+5)v adding the two= (v+5)^2 v^2 v^2 + 25 + 10v + v^2 ------- + ----- = -------------- ------ v(v+5) v(v+5) v(v+5) v(v+5) =2{v}^2+ 25 + 10v --------------- v(v+5)